OpenStudy (anonymous):

1. Kevin ran 4 miles more than Steve ran. The sum of their distances is 26 miles. How far did Steve run? The domain of the solution is {0, 4, 11, 22}. a. Define your variable. b. Write algebraic expressions for the distances each person ran. c. Write an equation that can help you solve the problem. d. Solve the equation. e. Check your answer.

OpenStudy (anonymous):

@jacob4900

OpenStudy (shadowlegendx):

Are you there?

OpenStudy (shadowlegendx):

Hello?

OpenStudy (shadowlegendx):

Ok it seems that you are afk so I will just solve the problem out for you :)

OpenStudy (anonymous):

lol sorry

OpenStudy (anonymous):

i'm just checking my answer is it 11

OpenStudy (shadowlegendx):

sum = the result of addition The sum of the miles Kevin and Steven ran is 26 miles. Turn this into an equation and you get K + S = 26 But we need to get only one variable defined here to be able to solve the equation. Remember what they said at the beginning of the problem. "Kevin ran 4 miles more than Steve ran." Make this into an equation and we get S + 4 = k So we can put S + 4 in for k and we will have only one variable defined So lets put our new data into our original equation of K + S = 26 New Equation: S + 4 + S = 26 Lets rewrite it for more better clarity: S + S + 4 = 26 Now lets solve: Like terms first: 2s + 4 = 26 Minus 4 from both sides and we get: 2s = 22 Divide both sides by 2 to isolate the variable and we get: s = 11 So now we figured out that Steve ran 11miles By knowing this we can figure out Kevin ran 15 miles as Kevin ran 4 more miles than Steve 11 + 4 = 15 Let me know if you need anymore help c;

OpenStudy (anonymous):

thank you so much :)

OpenStudy (shadowlegendx):

No problem :P